# Promptus: Can Prompts Streaming Replace Video Streaming with Stable Diffusion

### v1

@@ 3 Video to Prompt Inversion
This section describes how Promptus inverts frames into prompts. The workflow is illustrated in Figure 3. First, to ensure pixel alignment of the generated frames, a gradient descent-based prompt fitting framework is proposed (§3.1). Second, to control the bitrate of each prompt, a low-rank decomposition-based compression algorithm is introduced (§3.2). Third, to perform inter-frame compression on prompts, a temporally smooth prompt space is proposed (§3.3).
@@ 3.1 Gradient Descent based Prompt Fitting
Gradient Descent Framework. To obtain the inverse prompt, the most straightforward approach is to train a neural network to map the target image to a prompt. However, on one hand, this approach makes it difficult to ensure pixel alignment (like the aforementioned Captioning algorithm [46, 23]). On the other hand, as the inverse process of Stable Diffusion, this neural network needs to learn comparable knowledge, but this is very expensive (e.g., training a Stable Diffusion will cost between 600,000 and 10 million US dollars [12]). Therefore, instead of training a new neural network, we propose fully leveraging Stable Diffusion’s knowledge to infer the prompt. To this end, we adopt gradient descent to iteratively fit the prompt, with the framework shown in Figure 3. Specifically, at the beginning, the prompt is randomly initialized. Then, Stable Diffusion generates a frame based on this prompt. Since the prompt is random, the generated frame is meaningless. Third, the pixel-wise difference between the generated frame and the target frame is calculated as the loss value. Fourth, backpropagation is used to compute the gradient of the loss value with respect to the prompt. Finally, the prompt is updated using gradient descent. The above steps are iteratively executed until the loss value is sufficiently small, and the resulting prompt can satisfy pixel-aligned generation. In the above steps, Stable Diffusion is pre-trained and frozen, so it has prior knowledge. This knowledge is gradually distilled into the prompt through gradient descent fitting.
@@
Single-step denoising to avoid higher-order derivatives. As described in §2.2, Stable Diffusion uses iterative denoising to generate images. Therefore, the prompt recursively affects the generated image. This causes the gradient of the loss value with respect to the prompt to involve the computation of higher-order derivatives (such as 20th order), which introduces prohibitive computational and memory overhead. Therefore, instead of using the traditional Stable Diffusion, we adopt SD Turbo [9], a variant that can generate frames through single-step denoising. The adoption of SD Turbo allows gradient descent to only compute the first-order derivatives of the prompt, greatly improving efficiency. Although the quality of the generated frames is slightly weaker than the traditional iterative denoising Stable Diffusion, these quality losses can be compensated for in an end-to-end manner during prompt fitting.
@@
Employing embeddings as prompts to avoid non-differentiability. Computing the gradient of the prompt through backpropagation requires the forward computation from the prompt to the loss value to be completely differentiable. However, as shown in Figure 2, the forward computation includes the CLIP module [36], which converts text from discrete natural language to continuous embeddings. This step involves indexing and table lookup, making it non-differentiable. Gradients cannot be propagated to the text, preventing gradient descent. To address this, we discard the non-differentiable CLIP module and directly use embeddings as prompts for conditioning Stable Diffusion. In this case, the forward computation is fully differentiable, allowing gradient descent to be performed on the embeddings. In the following sections, prompt refers to the embedding rather than the text.
@@
Using a noisy previous frame instead of random noise to reduce latent space distance. According to §2.2, in addition to the prompt, the input random noise also affects the frames generated by Stable Diffusion. With the same prompt, different input noise results in different generated frames. At a high level, the input noise can be viewed as a point in the latent space, and the denoising process of Stable Diffusion actually moves this point under the control of the prompt. Therefore, the goal of Promptus is to find the inverse prompt that can move the point represented by the noise to the point of the target image in the latent space. Since we adopt a single-step denoising Stable Diffusion, if the random noise is far from the target image in the latent space, this movement cannot be completed in a single step, making it impossible to fit the inverse prompt. As shown in Figure 4(a), using random noise as input, after the loss value converges, the generated frame still has noticeable differences from the target frame, including blurring, noise artifacts, and inconsistent details. Therefore, we need to reduce the distance between the input noise and the target image in the latent space. We observe that in a video, adjacent frames are close in the latent space. Thus, we manually add noise to the previous frame as follows:
@@
Where $Z^{t-1}$ is the previous frame in the latent space. $N^{0}$ is a fixed noise. $\gamma$ is a hyperparameter that controls the degree of noise addition, which we set to $0.95$ in the experiment. $N^{t}$ is used as the noise input to Stable Diffusion for denoising the current frame. As shown in Figure 4(c), compared to random noise, the noisy previous frame can reduce the distance in the latent space, resulting in a generated frame that better matches the target frame.
Combining reconstruction and perceptual loss functions. To achieve pixel-aligned supervision, the most intuitive loss function is the per-pixel reconstruction loss, such as MSE. These reconstruction losses attempt to minimize the error of each pixel, while errors in high-frequency details and edges often lead to large pixel errors. Therefore, to reduce the overall error, reconstruction losses tend to abandon the fitting of edges and details, resulting in overly smooth and blurry images, as shown in Figure 4(b). To make the generated frames sharp and clear, one approach is to use perceptual loss instead of reconstruction loss, such as LPIPS [55]. Perceptual loss is based on deep learning and can estimate the subjective quality of images as perceived by the human eye. Since the human eye is highly sensitive to image details, perceptual loss can make the generated images sharper with richer details. However, perceptual loss aims to maximize the overall subjective quality of the image without focusing on the exact consistency of each pixel, leading to misalignment between the generated and target images. Therefore, to simultaneously ensure pixel alignment and subjective quality, we combine the reconstruction and perceptual loss as the fitting loss $D$:
@@
where $D_{rec}$ represents the reconstruction loss, which is MSE by default. $D_{per}$ represents the perceptual loss, which is LPIPS by default. $\alpha$ is a hyperparameter that balances pixel alignment and perceptual quality. In our experiments, we set $\alpha$ to 0.8. The final result is shown in Figure 4(c). It can be observed that by jointly optimizing the reconstruction loss and the perceptual loss, the images generated by Promptus ensure pixel-level alignment while maintaining a sharp appearance, making them almost identical to the ground truth images.
@@
3.2 Low-rank Decomposition based Prompt Bitrate Control
@@
According to §3.1, Promptus uses embeddings as prompts instead of text. However, embeddings are $m*n$ matrices (e.g., $1024*77$), where each element is a high-bit floating-point number (e.g., 32-bit float type), resulting in a much higher bitrate. To reduce the bitrate of prompts, Promptus has two directions: First, dimensionality reduction decreases the number of parameters in the prompt. Second, quantization reduces the number of bits for each parameter in the prompt.
@@
Low-rank matrix decomposition. To perform dimensionality reduction on the prompt, the most straightforward method is to first fit the complete prompt and then perform dimensionality reduction algorithms such as SVD (Singular Value Decomposition) or PCA (Principal Component Analysis) on it. However, these methods only perform dimensionality reduction based on the data distribution of the prompt, ignoring the impact of prompt degradation on the generation results of Stable Diffusion. This inevitably leads to a degradation in the quality of the generated images. Therefore, instead of performing explicit dimensionality reduction, Promptus proposes to directly fit a low-dimensional prompt end-to-end, thereby reducing the quality degradation caused by dimensionality reduction. To achieve this, Promptus integrates the inverse process of CANDECOMP/PARAFAC decomposition [22] into gradient descent fitting. Specifically, Promptus calculates the embedding $c$ as follows:
$\vspace{-6mm}c=\frac{u*v}{\sqrt{r}}$ (3)
@@
where $u$ and $v$ are two low-rank factor matrices, and $r$ is the rank of the embedding. $u$ and $v$ compose the embedding $c$ through outer product and normalization. At this point, the embedding $c$, as an intermediate variable, is no longer fitted or stored. $u$ and $v$, as the new representation of the prompt, will be randomly initialized and fitted.
@@
The rank $r$ determines the trade-off between bitrate and quality. Compared to the embedding size of $m*n$ (e.g., $1024*77$), the total size of $u$ and $v$ is $(m+n)*r$. Therefore, reducing $r$ significantly lowers the bitrate. However, on the other hand, when Rank $r$ is smaller, the embedding is constrained to be a low-rank matrix, resulting in weaker representational capability and inability to fit high-frequency details in the image, as shown in Figure 7. Consequently, it is necessary to dynamically select the most appropriate $r$ based on the currently available network bandwidth to trade off between bitrate and quality.
@@
Fitting-aware quantization. Although the number of parameters in the prompt has been significantly reduced through low-rank matrix decomposition, each parameter in $u$ and $v$ is still a high-bit floating-point number (such as 32-bit float type). Therefore, to further reduce the bitrate, it is necessary to quantize $u$ and $v$, reducing the number of bits for each parameter (such as lowering it to 8 bits). The most straightforward approach is to first fit $u$ and $v$ and then perform quantization. However, this post-quantization technique inevitably leads to quality loss. To address this, Promptus incorporates quantization into the fitting process, automatically compensating for the quantization loss through end-to-end gradient descent. We tested the impact of different quantization configurations on quality. Compared to traditional post-quantization, our method can reduce the number of bits from 32 to 8 with almost no quality loss.
@@
3.3 Prompt inter-frame compression based on temporal smoothing
Figure 5: At the receiver, Promptus approximates the unreceived frames by linear interpolation in the prompt space.
According to §3.1, Promptus fits each frame of the video as an independent prompt, without considering the correlation between prompts across frames. Therefore, during video streaming, Promptus needs to transmit a prompt for each frame, resulting in a linear increase in bitrate as the frame rate rises. However, codec-based streaming can avoid this problem through inter-frame compression. Is it possible to perform inter-frame compression on prompts as well? The most straightforward solution is to reshape the prompt of each frame into a two-dimensional matrix and encode it using a video codec. However, unlike conventional videos, we found that this reshaped prompt looks like random noise, lacking any patterns, structures, or smooth regions, causing video codecs to no longer work.
For inter-frame compression of prompts, our insight is: prompts are high-level semantics, so the prompts of continuous video frames should change continuously. If two temporally close frames are also sufficiently close in the prompt space, then the prompts of the frames between these two frames can be approximated by linear interpolation. With this, during streaming, we only need to sparsely transmit the prompts of a few frames (as keyframes), and the prompts of the remaining frames can be obtained by linear interpolation of the keyframe prompts, as illustrated in Figure 5. Since only a small portion of keyframes need to transmit prompts, inter-frame compression is achieved.
To ensure that adjacent frames are sufficiently close in the prompt space, we add temporal smoothing regularization to the embedding during fitting, as follows:
$\lambda=\left\|c^{t}-c^{t-1}\right\|_{2}$ (4)
Here, $c^{t}$ represents the embedding of the frame currently being fitted, while $c^{t-1}$ denotes the embedding of the previously fitted frame. The final loss function $L$ is as follows:
$L=\beta*D+(1-\beta)*\lambda$ (5)
Where $\beta$ is a hyperparameter used to balance the fitting loss and the temporal smoothing regularization. In our experiments, we set it to 0.2.
@@ 3.3 Prompt inter-frame compression based on temporal smoothing
Figure 6: Visualization of the prompt interpolation results. When applying temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
Temporal smoothing regularization works, as shown in Figure 6. Without temporal smoothing regularization, the interpolation results suffer from severe distortions, disrupting the video’s content and structure. With temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
4 Prompt Streaming System Design and Implementation
@@
§3 shows how Promptus inverts videos into pixel-aligned prompts. In this section, we will further describe how Promptus utilizes this Prompt Inversion technique to design and implement a streaming system. Specifically, it will be divided into two parts: the sender side and the receiver side.

### v2

@@ 3 Video to Prompt Inversion
This section describes how Promptus inverts frames into prompts. The workflow is illustrated in Figure 3. First, to ensure pixel alignment of the generated frames, a gradient descent-based prompt fitting framework is proposed (§3.1). Second, to control the bitrate of each prompt, a low-rank decomposition-based adaptive bitrate algorithm is introduced (§3.2). Third, to perform inter-frame compression on prompts, an interpolation-aware fitting algorithm is proposed (§3.3).
@@ 3.1 Gradient Descent based Prompt Fitting
Gradient Descent Framework. In order to obtain the inverse prompt, the most straightforward approach is to train a neural network to map the target image to a prompt. However, on one hand, this approach makes it difficult to ensure pixel alignment (like the aforementioned Captioning (yang2023vid2seq, ; hu2022scaling, )). On the other hand, as the inverse process of Stable Diffusion, this neural network needs to learn comparable knowledge, but this is very expensive (e.g., training a Stable Diffusion will cost between 600,000 and 10 million US dollars (SDcost, )). Instead of training a new neural network, we propose fully leveraging Stable Diffusion’s knowledge to infer the prompt. To this end, we adopt gradient descent to iteratively fit the prompt, with the framework shown in Figure 3. Specifically, at the beginning, the prompt is randomly initialized. Stable Diffusion generates a frame based on this prompt. Since the prompt is random, the generated frame is meaningless. Third, the pixel-wise difference between the generated frame and the target frame is calculated as the loss value. Fourth, backpropagation is used to compute the gradient of the loss with respect to the prompt. Finally, the prompt is updated using gradient descent. The above steps are iteratively executed until the loss is sufficiently small, and the resulting prompt can satisfy pixel-aligned generation. In the above steps, Stable Diffusion is pre-trained and frozen, so it has prior knowledge. This knowledge is distilled into the prompt via gradient descent.
@@
Single-step denoising to avoid unrolled long gradient chains. As described in §2.2, Stable Diffusion uses iterative denoising to generate images. So the prompt recursively influences the generated image. This causes the gradient of the loss value with respect to the prompt to involve unrolling numerous iterations (e.g., 50 steps), where the accumulated long gradient chain introduces prohibitive computational and memory overhead. Thus, instead of using the traditional Stable Diffusion, we adopt SD (2.1) Turbo (sdturbo, ), a variant that can generate frames through single-step denoising. The adoption of SD Turbo allows gradient descent to require only a single forward-backward pass, greatly improving efficiency. Although the quality of the generated frames is weaker than the traditional Stable Diffusion, these quality losses can be compensated for in an end-to-end manner during fitting.
@@
Employing embeddings as prompts to prevent non-differentiability. Computing the gradient of the prompt through backpropagation requires the forward computation from the prompt to the loss value to be completely differentiable. However, as described in §2.2, the forward computation includes the CLIP module (radford2021learning, ), which converts text from discrete natural language to continuous text embeddings. This conversion involves indexing and table lookup, making it non-differentiable. Gradients cannot be propagated to the text in natural language, preventing gradient descent. To address this, we discard the non-differentiable CLIP module and directly use text embeddings as prompts for conditioning Stable Diffusion. In this case, the forward computation is fully differentiable, allowing gradient descent to be performed on the text embeddings. In the following sections, prompt refers to the text in embedding rather than the text in natural language [^1] It is important to note that, even though it is not natural language, embeddings can still serve as prompts. On one hand, prompts can take multiple modalities, such as images (zhang2023adding, ; blattmann2023stable, ), audio (biner2024sonicdiffusion, ; tang2024any, ), embeddings (gal2022image, ; ruiz2023dreambooth, ; kawar2023imagic, ; pan2022extreme, ), fMRI (chen2024cinematic, ; takagi2023high, ), etc. On the other hand, the fundamental characteristic of prompts is that they serve as conditions to guide generation. The resolution of the generated videos is only dependent on the size of the random noise, unrelated to the prompt size, as described in §2.2. So prompts have the advantage of being decoupled from video resolution compared to encoding..
@@
Using a noisy previous frame instead of random noise to reduce latent space distance. According to §2.2, in addition to the prompt, the input random noise also affects the generated frames. With the same prompt, different input noise results in different generated frames. At a high level, the input noise can be viewed as a point in the latent space, and the denoising process actually moves this point under the control of the prompt. Therefore, the goal of Promptus is to find the inverse prompt that can move the point represented by the noise to the point of the target image in the latent space. Since we adopt a single-step denoising Stable Diffusion, if the input noise is far from the target image in the latent space, this movement cannot be completed in a single step, making it impossible to fit the inverse prompt. As shown in Figure 4(a), using random noise as input, after the loss value converges, the generated frame still has noticeable differences from the target frame, including blurring, noise artifacts, and inconsistent details. Therefore, we need to reduce the distance between the input noise and the target frame in the latent space. We observe that in a video, adjacent frames are close in the latent space. Thus, we manually add noise to the previous frame as follows:
@@
Where $Z^{t-1}$ is the previous frame in the latent space. $N^{0}$ is a fixed noise. $\gamma$ is a hyperparameter that controls the degree of noise addition, which we set to $0.95$ in the experiment. $N^{t}$ is used as the noise input to Stable Diffusion for generating the current frame. As shown in Figure 4(c), compared to random noise, the noisy previous frame can reduce the distance in the latent space, resulting in a generated frame that better matches the target frame.
Combining reconstruction and perceptual loss functions. To achieve pixel-aligned supervision, the most intuitive loss function is the per-pixel reconstruction loss, such as MSE. These reconstruction losses attempt to minimize the error of each pixel, while errors in high-frequency details and edges often lead to large pixel errors. Therefore, to reduce the overall error, reconstruction losses tend to abandon the fitting of edges and details, resulting in overly smooth and blurry images, as shown in Figure 4(b). To make the generated frames sharp and clear, one approach is to use perceptual loss instead of reconstruction loss, such as LPIPS (zhang2018unreasonable, ). Perceptual loss is based on deep learning and can estimate the subjective quality of images as perceived by the human eye. Since the human eye is highly sensitive to image details, perceptual loss can make the generated images sharper with richer details. However, perceptual loss aims to maximize the overall subjective quality of the image without focusing on the exact consistency of each pixel, leading to misalignment between the generated and target images. To simultaneously ensure pixel alignment and subjective quality, we combine the reconstruction and perceptual loss as the fitting loss $D$:
@@
where $D_{rec}$ represents the reconstruction loss, which is MSE by default. $D_{per}$ represents the perceptual loss, which is LPIPS by default. $\alpha$ is a hyperparameter that balances pixel alignment and perceptual quality. In our settings, we set $\alpha$ to 0.8. The final result is shown in Figure 4(c). It can be observed that by jointly optimizing the reconstruction loss and the perceptual loss, the image generated by Promptus ensures pixel-level alignment while maintaining a sharp appearance, making them almost identical to the ground truth images.
Prompt Regularization. We found that even with real video as supervision, fitting occasionally fails. This is because as gradient descent progresses, the distribution of prompts gradually shifts away from the distribution of CLIP embeddings, resulting in an inability to generate meaningful images. To address this, we propose prompt regularization to constrain the distribution of prompts:
$\lambda=\left\|\overline{c}-\mu\right\|_{2}$ (3)
Here, $\mu$ is the mean of embeddings output by CLIP (which we obtained through statistics on a large amount of text, specifically $\mu=-0.168$). The final loss function $L$ is as follows:
$L=\beta*D+(1-\beta)*\lambda$ (4)
Where $\beta$ is a hyperparameter used to balance the fitting loss and the prompt regularization. In our experiments, we set it to 0.9.
@@
3.2 Low-rank Decomposition based Prompt Adaptive Bitrate Control
@@
According to §3.1, each prompt is an $m*n$ embedding matrix (e.g., $1024*77$), where each element is a floating-point number (e.g., 32-bit float), resulting in a fixed and high bitrate. To precisely control the bitrate of prompts, Promptus has two directions: First, dimensionality reduction decreases the number of prompt parameters. Second, quantization reduces the number of bits for each parameter.
@@
Low-rank matrix decomposition. To perform dimensionality reduction on the prompt, the most straightforward method is to first fit the complete prompt and then perform dimensionality reduction algorithms such as Singular Value Decomposition or Principal Component Analysis on it. However, these methods only perform dimensionality reduction based on the data distribution of the prompt, ignoring the impact of prompt degradation on the generation results. This inevitably leads to a degradation in the quality of the generated images. Therefore, instead of performing explicit dimensionality reduction, Promptus proposes to directly fit a low-dimensional prompt end-to-end, thereby reducing the quality degradation caused by dimensionality reduction. To achieve this, Promptus integrates the inverse process of CANDECOMP/PARAFAC decomposition (harshman1970foundations, ) into gradient descent fitting. Specifically, Promptus calculates the embedding $c$ as follows:
$c=\frac{u*v}{\sqrt{r}}$ (5)
@@
where $u$ and $v$ are two low-rank factor matrices with dimensions $m*r$ and $r*n$, respectively. $r$ is the rank of the embedding. $u$ and $v$ compose the embedding $c$ through outer product and normalization. At this point, the embedding $c$, as an intermediate variable, is no longer fitted or stored. $u$ and $v$, as the new representation of the prompt, will be randomly initialized and fitted.
@@
The rank $r$ determines the trade-off between bitrate and quality. Compared to the embedding size of $m*n$ (e.g., $1024*77$), the total size of $u$ and $v$ is $(m+n)*r$. So reducing $r$ significantly lowers the bitrate. However, on the other hand, when Rank $r$ is smaller, the embedding is constrained to be a low-rank matrix, resulting in weaker representational capability and inability to fit high-frequency details in the image, as shown in Figure 7. So it is necessary to dynamically select the most appropriate $r$ based on the current network bandwidth to trade off between bitrate and quality.
@@
Fitting-aware quantization. Although the number of parameters in the prompt has been significantly reduced through low-rank matrix decomposition, each parameter in $u$ and $v$ is still a high-bit floating-point number (such as 32-bit float type). Therefore, to further reduce the bitrate, it is necessary to quantize $u$ and $v$, reducing the number of bits for each parameter. We adopt differentiable fake quantization (zafrir2019q8bert, ) and incorporate quantization into the fitting process, automatically compensating for the quantization loss through end-to-end gradient descent. We test the impact of different quantization configurations on quality. The results indicate our method reduces the number of bits from 32 to 8 without quality loss.
@@
3.3 Interpolation-aware Fitting based Prompt Inter-frame Compression
@@
Figure 5: The prompt interpolation not only fully preserves the details but also successfully fit the motion.
Figure 6: At the receiver, Promptus approximates the unreceived frames by linear interpolation in the prompt space. Thus, instead of sending the prompt for each frame, only the prompts for keyframes need to be transmitted sparsely, thereby achieving inter-frame compression.
According to §3.1, Promptus fits each frame of the video as an independent prompt, without considering the correlation across frames. Therefore, during video streaming, Promptus needs to transmit a prompt for each frame, resulting in a linear increase in bitrate as the frame rate rises. However, codec-based streaming can avoid this problem through inter-frame compression. Is it possible to perform inter-frame compression on prompts as well? The most straightforward solution is to reshape the prompt of each frame into a two-dimensional matrix and encode it using a video codec. However, unlike conventional videos, we found that this reshaped prompt looks like random noise, lacking any patterns, structures, or smooth regions, causing codecs to no longer work.
For inter-frame compression of prompts, our insight is: prompts are high-level semantics, so the prompts of continuous video frames can change continuously. With this, during streaming, we only need to sparsely transmit the prompts of a few frames (as keyframes), and the prompts of the remaining frames can be obtained by linear interpolation of the keyframe prompts, as illustrated in Figure 6. Since keyframes account for only a small portion of the total frames, inter-frame compression is achieved.
To ensure that the interpolated prompts can generate meaningful frames, we integrate prompt interpolation into the gradient descent fitting. Specifically, we randomly initialize prompts for keyframes, interpolate prompts for each intermediate frame, and generate frames. At this point, the prompts for intermediate frames exist as intermediate variables, so they are not optimized; instead, gradients are backpropagated to keyframes to directly optimize their prompts.
Prompt interpolation works, as shown in Figure 5, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video. In practice, scene changes often occur within the video, at which point prompt interpolation no longer works. Therefore, we will continuously detect scene changes and treat the new scenes as new videos. For more details, please refer to §4.1.
4 Prompt Streaming System Implementation
@@
§3 shows how Promptus inverts videos into pixel-aligned prompts. In this section, we will further describe how Promptus utilizes this Prompt Inversion technique to design a streaming system. Specifically, it will be divided into two parts: the sender side and the receiver side.